OpenStudy (jumperman):

How do you factor trinomials? Here is an example of one: x2 - 8x - 20

OpenStudy (anonymous):

http://ts4.mm.bing.net/th?id=HN.608013394737893209&w=257&h=179&c=7&rs=1&pid=1.7 You want me to walk you through it?

OpenStudy (jumperman):

Please.

OpenStudy (precal):

@nelsonjedi FOIL is used to expand (x+a)(a+b) out not to factor

OpenStudy (jumperman):

2 and 8?

OpenStudy (jumperman):

Well how do you factor?

OpenStudy (precal):

take the last term 20

OpenStudy (precal):

list all the factors of 20

OpenStudy (precal):

1 times 20 2 times 10 4 times 5

OpenStudy (anonymous):

My bad.

OpenStudy (precal):

@nelsonjedi no problem

OpenStudy (precal):

the minus sign in front of 20 ie -20 tells you to subtract the factors

OpenStudy (anonymous):

I use FOIL as an example and then tell them to work backwards

OpenStudy (jumperman):

So 1, 2, 4, 5, 10, and 20 are all the factors?

OpenStudy (precal):

20-1 =19 10-2=8 5-4=1 use the ones that will give you positive numbers so don't use 1-20 =-19 we don't care about the signs at this point, we do the signs separately

OpenStudy (precal):

we are looking for the middle term which is 8 so which factors do we use?

OpenStudy (jumperman):

10-2.

OpenStudy (precal):

by the way if the middle term is not listed then the trinomial is what is known as "prime" and can not be factored but in this case we can find 8

OpenStudy (precal):

ok so let's break it up

OpenStudy (precal):

x times x is x to the second power, agreed?

OpenStudy (jumperman):

Yes.

OpenStudy (precal):

2 times 10 will give you 20